OpenStudy (anonymous):

Let u be a constant. Prove that u=sin(x-at)+ln(x+at) is a solution to the wave equation (u_tt=a^2*u_xx)..?

OpenStudy (turingtest):

plug it in and differentiate

OpenStudy (turingtest):

what are \(u_{xx}\) and \(u_{tt}\) ?

OpenStudy (anonymous):

Oh, hahaha I was looking at this way wrong. So what I have to do are take the second partial derivatives of it and just plug them in, right?

OpenStudy (turingtest):

yup. it's easy to over-think problems when they're worded all fancy I know :P

OpenStudy (anonymous):

Hahaha thank you for all of your help, it is greatly appreciated :)

OpenStudy (anonymous):

So, I got f_x=2x+ky f_y=2y+kx. I then did this: f_x=2x+ky=0 f_y=2y+kx=0 I solved for y and substituted y in the f_x equation with the value I got. I then had: y=kx/2 2x+ky=0------>2x+k(kx/2) I then proceeded to solve for x and got... x(4+k^2)=0 So...x=0? I then plugged 0 into the y= equation y=k(0)/2 and got y=0. Hence (0,0) is a critical point, yes? What's my next step?

OpenStudy (anonymous):

SORRY, WRONG QUESTION